The indicator methyl red has a pKHIn = 4.95. It changes from red to yellow over the pH range from 4.4 to 6.2.

If the indicator is placed in a colourless buffer solution of pH = 4.55, what percent of the indicator will be present in the acid form, HIn?


1
Expert's answer
2021-07-05T03:24:35-0400

K = 10-4.95 = 1.122 ∙10-5

1.122 ∙10-5 = [A-][H+] / [HA]

[H+] = 10-4.55 = 2.818 ∙10-5

[HA] = 1-[A-]

1.122 ∙10-5 = [A-] ∙ (2.818 ∙10-5) / (1-[A-])

0.398 = [A-] / (1 - [A-])

[A-] = 0.28

[HA] = 1 - 0.28 = 0.72

C%(HA) = 0.72